# Skewed score: a statistical framework to assess autograders

### v1

@@ 1.2 Integrating the autograder assessment with the research question
Recent literature has raised concerns that autograders may demonstrate self-bias, a tendency to assign better scores to outputs generated by the same base model [Panickssery et al., 2024; Liu et al., 2024b; Koo et al., 2024] or outputs from models vs humans [Liu et al., 2023], which would naturally compromise evaluation automation efforts. Similarly to above, with a GLM framework, we can seamlessly integrate this analysis into a specific research question and quantify self-bias for a particular application.
@@
To answer these questions, she needs repeated responses for the same items. Until now, we have assumed that each data point corresponds to a different item, so lets instead imagine that Florence’s dataset consists of four items (for simplicity), with each model answering each item 25 times. This results in 50 responses per item (25 per model), and each grader scores all of the attempts. The data split by items can be seen in the left panel of fig. 5 (different items are represented by violin plots of the same colour).
To analyse this, we extend eq. 3 by including two additional terms. First, a main effect for items, to capture whether some items receive systematically higher or lower scores. Second, a grader–item interaction, to test whether certain graders behave differently on specific items.
@@
Here, $\phi_{i}$ is the linear predictor on the latent scale, and $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into ordered categories. The term $\beta_{1}(X_{i}^{\text{grader}})$ captures the main effect of each grader, and $\beta_{2}$ models the effect of LLM identity (e.g., whether the response was produced by LLM A or B). The new term $\beta_{3}(X_{i}^{\text{item}})$ represents the main effect of each item which captures whether some questions tend to receive higher or lower scores overall. The final term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, captures grader–item interactions and is implemented in the same way as the interaction term in eq. 4: rather than a single multiplicative term, this is a set of coefficients (one for each grader–item combination) estimated using dummy coding without a sum-to-zero constraint. This allows us to directly compare individual combinations and detect whether certain graders are more lenient or harsh on specific items. As before, all main categorical effects (grader, item) are encoded using effect coding, so that the resulting coefficients reflect deviations from the overall mean.
After fitting the model, Florence inspects the estimated effects (right panel of fig. 5). She looks at two things:
@@
    The main effect of item ($\beta_{3}$ in eq. 6). She sees that some items receive consistently higher scores than others, e.g., Item 1 has a notably positive main effect, suggesting that graders give higher scores for these responses. In contrast, Item 4 has a negative effect, suggesting that graders give lower scores on this item. Overall this suggests that item 1 is easier to answer, and item 4 harder to answer.
@@
    The grader–item interaction term ($\beta_{4}$ in eq. 6) to see whether graders disagree more on particular questions. Most interaction estimates are small and their credible intervals overlap with zero, suggesting no strong evidence that specific graders behave differently on individual items.
From this, Florence concludes that while some items appear easier than others, grader disagreement does not seem to be concentrated on any particular question. She might also be interested in quantifying the overall level of agreement between graders. A common approach in such cases is to compute an inter-rater agreement statistic, such as Krippendorff’s $\alpha$ [Bavaresco et al., 2025; Tam et al., 2024], which measures the consistency of annotations across (more than 2) raters.
@@
However, while helpful, such agreement metrics have important limitations: they do not distinguish between disagreement due to noise and disagreement due to systematic bias. In fig. 6, both the traditional and GLM-based $\alpha$ estimates are below zero, suggesting substantial disagreement. But this does not tell us why graders disagree. To answer that, Florence turns back to the right panel of fig. 5. There, she observes that human graders have a strong positive main effect relative to autograders, meaning they consistently assign higher scores. She also notes that while items differ in average difficulty, the grader–item interaction effects are close to zero, indicating that graders do not differ much in how they score specific items.
Together, this pattern suggests that the low $\alpha$ is not due to random noise, but to a systematic scoring bias across grader types. To illustrate this, Florence goes one step further: for each predicted score, she subtracts the estimated main effect of the corresponding grader (i.e., their bias term from the model) from the linear predictor before mapping it to a categorical score. This yields a set of bias-adjusted predictions that reflect how graders might score if they did not differ systematically in their baseline scoring. In essence, Florence is constructing a counterfactual scenario: “What would agreement look like if graders did not have systematic differences?”
@@
In conclusion, this framework allows Florence to go beyond traditional agreement metrics and understand the source of disagreement. She can still compute familiar statistics such as Krippendorff’s $\alpha$, Cohen’s $\kappa$, or Kendall’s $\tau$, but do so using posterior predictions from the GLM. These predictions account for covariates in the data and yield credible intervals for each metric, enabling more robust and interpretable comparisons across grader types, task formats, and evaluation settings.
@@
So far, we have focused on evaluation settings where graders assign absolute scores. However, many LLM evaluations instead rely on pairwise comparisons, where graders are asked to choose which of two outputs better satisfies a target criterion (e.g., correctness). The same statistical modeling framework can be applied in this setting, with the outcome modeled as a binary preference. We use such a setup to illustrate how pairwise comparisons can be modeled and to examine length bias, which has often been observed in pairwise evaluation settings. Of course, length bias can also be captured in absolute score setups similarly to other biases in previous sections.
Let’s imagine that Florence is interested in comparing the outputs of 3 LLMs. In a pairwise setting, she evaluates outputs from two different models at a time. To model this, we use a binomial GLM with a logistic link function. The outcome variable $y_{i}$ indicates whether the first model in the pair was chosen ($y_{i}=1$) or not ($y_{i}=0$), and we include a categorical effect in the model to specify the specific LLM pair being compared.
If Florence has two (or more) autograders selecting their favourite from each pair, she can, as in previous questions, add a parameter for the autograder which will capture specific autograder behaviour. For simplicity we omit this here.
@@
Florence’s younger brother, always up-to-date with ML controversies, recently told her that some autograders may systematically prefer longer outputs even when those outputs are not of higher quality, a phenomenon commonly referred to as length bias [Zheng et al., 2023; Dubois et al., 2024]. To test this formally, she adds a continuous predictor capturing the token-length difference between the two outputs to the model, and compares the model with and without this term (cf. Appendix). For demonstration purposes, let’s look at the model with length bias here:
@@
$\begin{gathered}y_{i}\sim\text{Binomial}(1,p_{i})\\ \text{logit}(p_{i})=\beta_{0}+\beta_{1}(\text{LLM}_{i}^{(1)},\text{LLM}_{i}^{(2)})+\beta_{2}\cdot\text{lengthDiff}_{i}\end{gathered}$ (7)
In this model, $y_{i}$ is a binary outcome indicating whether the first LLM in the comparison was preferred. The model uses a logistic link function to predict the log-odds of selecting the first LLM, based on three components. The intercept $\beta_{0}$ captures the overall tendency to prefer the first LLM, averaged across all graders and model pairs. The term $\beta_{1}(\text{LLM}_{i}^{(1)},\text{LLM}_{i}^{(2)})$ encodes the contrast associated with the specific ordered pair of LLMs. With three models in the pool (A, B, and C), there are three combinations (A-B, A-C and B-C) and corresponding parameters. The predictor $\text{lengthDiff}_{i}$ is the token-length difference between the two outputs. The coefficient $\beta_{2}$ reflects how much this length difference affects the probability of selecting the first LLM. A positive $\beta_{2}$ suggests that graders systematically favour longer outputs, regardless of quality.
@@
Importantly, once we have estimated the probability of choosing one LLM over another, we can compare these probabilities across pairs. This allows us to identify both rational (transitive) and irrational (intransitive) patterns of decision-making, such as cyclic dependencies (e.g., preferring A over B, B over C, but C over A). Such intransitivities have been shown to exist in LLM evaluations [Xu et al., 2025]. Traditional models like the Bradley–Terry model implicitly assume transitivity and thus cannot capture these cycles. Recent approaches have proposed either removing intransitivities from datasets [Yu et al., 2025] or explicitly quantifying them [Zhang et al., 2025; Liu et al., 2024a; Zhao et al., 2024]. A notable strength of our approach is that it naturally captures these intransitivities alongside grader biases and differences in LLM performance, eliminating the need for supplementary analyses.
@@
Inspecting the estimated effects (cf. Figure), we observe a positive effect for $\beta_{2}$. From this, Florence can infer that the autograder is more likely to select longer outputs, regardless of their actual quality. In other words, there appears to be an implicit association between output length and perceived correctness. By quantifying such biases in the model, Florence can account for different effects when interpreting LLM rankings. For example, if LLM A wins most comparisons but also consistently produces longer outputs, she might ask: "Is LLM A genuinely better, or just longer? Can I really trust my autograder?" This integrated framework allows her to disentangle such effects and draw more thorough conclusions.
@@ Conclusion
In this paper, we introduced a statistical framework for evaluating autograders using Bayesian generalised linear models. Importantly, by jointly modelling the evaluation outcome and the scoring process, this approach enables researchers to assess both LLM performance and grader behaviour within a single, interpretable structure.
Through a series of examples, we followed a fictional character, Florence, and her journey toward evaluating autograders. We used simulated data throughout to explore a wide range of setups and to highlight key modelling principles in a controlled and reproducible way.
In particular, we showed how this framework can be used to compute more precise estimates of inter-rater agreement and, in contrast to traditional metrics, to understand the sources of disagreement. We demonstrated how it can quantify various types of biases (e.g., self-bias and length bias), how to capture individual-level differences (both for individual graders and individual items) and how to improve the estimation of group-level trends through hierarchical modelling (especially useful in settings with limited or imbalanced data). We also demonstrated the flexibility of this method by showing how it can easily be adapted to different evaluation formats (i.e., absolute scoring and pairwise comparisons).
The examples presented are by no means exhaustive. Many other applications and extensions of GLMs are possible, and we hope this work provides a clear and practical starting point for researchers seeking to adapt the framework to their own evaluation scenarios. To support practical adoption, we have summarised common evaluation questions and their implementation in a GLM setup in table 1. All models presented in this paper are implemented in the open-source HiBayes package. We additionally provide reproducible notebooks for the examples here.
We hope that this approach will serve as a practical and extensible tool for understanding autograder behaviour and strengthening the reliability of LLM evaluation across diverse tasks and settings.
@@ Appendix C Supplementary question: Is the grading scale well calibrated?
Having an ordinal scale, instead of just a description, is of course very useful. But this scale is not as simple to interpret as an interval scale (e.g., temperature). If Florence wants to make claims like "autograder A increases the score by 1 point," she needs to be aware of whether the intervals between categories are equivalent. If they are not (which is often the case with ordinal scales), a 1-point increase will mean something different depending on where it occurs on the scale. For example, moving from 5 to 6 might represent crossing some fundamental threshold, while moving from 9 to 10 might just be a small qualitative improvement between two already good responses.

### v2

@@ 1.2 Integrating the autograder assessment with the research question
Recent literature has raised concerns that autograders may demonstrate self-bias, a tendency to assign better scores to outputs generated by the same base model [Panickssery et al., 2024; Liu et al., 2024b; Koo et al., 2024] or outputs from models vs. humans [Liu et al., 2023], which would naturally compromise evaluation automation efforts. Similarly to above, with a GLM framework, we can seamlessly integrate this analysis into a specific research question and quantify self-bias for a particular application.
@@
To answer these questions, she needs repeated responses for the same items. Until now, we have assumed that each data point corresponds to a different item, so lets instead imagine that Florence’s dataset consists of four items (for simplicity), with each model answering each item 25 times. This results in 50 responses per item (25 per model); each grader scores all of the attempts. The data split by items can be seen in the left panel of fig. 5 (different items are represented by violin plots of the same colour).
To answer Question 4, we extend eq. 3 by including two additional terms. The first term, $\beta_{3}(X_{i}^{\text{item}})$, accounts for a main effect of items, capturing whether some items receive systematically higher or lower scores. The second term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, represents a grader–item interaction, allowing us to test whether particular graders behave differently on specific items.
@@
Here, $\phi_{i}$ is the linear predictor on the latent scale, and $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into ordered categories. The term $\beta_{1}(X_{i}^{\text{grader}})$ captures the main effect of each grader, and $\beta_{2}$ models the effect of LLM identity (e.g., whether the response was produced by LLM A or B). As mentioned above, the new term $\beta_{3}(X_{i}^{\text{item}})$ represents the main effect of each item, which captures whether some questions tend to receive higher or lower scores overall. The final term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, captures grader–item interactions and is implemented in the same way as the interaction term in eq. 4: rather than a single multiplicative term, this is a set of coefficients (one for each grader–item combination) estimated using dummy coding without a sum-to-zero constraint. This allows us to directly compare individual combinations and detect whether certain graders are more lenient or harsh on specific items. As before, all main categorical effects (grader, item) are encoded using effect coding, so that the resulting coefficients reflect deviations from the overall mean.
After fitting the model, Florence inspects the estimated effects (right panel of fig. 5). She focuses on two aspects:
@@
    The main effect of item ($\beta_{3}$ in eq. 6), observing that some items consistently receive higher or lower scores. For example, Item 1 shows a notably positive main effect, indicating that graders tend to assign higher scores to responses for this item. Conversely, Item 4 exhibits a negative effect, implying that graders give lower scores on this item. Overall, this suggests that Item 1 is easier to answer, while Item 4 is more challenging.
@@
    The grader–item interaction term ($\beta_{4}$ in eq. 6), which helps identify whether graders differ more on specific questions. Most interaction estimates are small and their credible intervals overlap with zero, indicating weak evidence that specific graders behave differently on individual items.
From these results, Florence concludes that while some items appear easier than others, grader disagreement is not concentrated on any particular question. However, she still wants to understand how consistently graders agree overall across all items. A common way to quantify this is to compute an inter-rater agreement statistic such as Krippendorff’s $\alpha$ [Tam et al., 2024; Bavaresco et al., 2025], which measures the consistency of annotations across (more than 2) raters.
@@
However, while helpful, such agreement metrics have important limitations: they do not distinguish between disagreement due to noise and disagreement due to systematic bias. In fig. 6, both the traditional and GLM-based $\alpha$ estimates are below zero, suggesting substantial disagreement. But this does not tell us why graders disagree. To answer this question, Florence turns back to the right panel of fig. 5, where she observes that human graders have a strong positive main effect relative to autograders, meaning they consistently assign higher scores. She also notes that while items differ in average difficulty, the grader–item interaction effects are close to zero, indicating that graders do not differ much in how they score specific items.
Together, these two observations suggest that the low $\alpha$ is not due to random noise, but to a systematic scoring bias across grader types. To illustrate this, Florence goes one step further: for each predicted score, she subtracts the estimated main effect of the corresponding grader (i.e., their bias term from the model) from the linear predictor before mapping it to a categorical score. This yields a set of bias-adjusted predictions that reflect how graders might score if they did not differ systematically in their baseline scoring. In essence, Florence is constructing a counterfactual scenario: “What would agreement look like in the absence of systematic differences among graders?”
@@
In conclusion, this framework allows Florence to go beyond traditional agreement metrics and understand the source of disagreement. She can still compute familiar statistics such as Krippendorff’s $\alpha$, Cohen’s $\kappa$, or Kendall’s $\tau$, but now using posterior predictions from the GLM. These predictions account for covariates in the data and yield credible intervals for each metric, enabling more robust and interpretable comparisons across grader types, task formats, and evaluation settings.
@@
So far, we have focused on evaluation settings where graders assign absolute scores. However, many LLM evaluations rely on pairwise comparisons, where graders are asked to choose which of two outputs better satisfies a target criterion (e.g., correctness). The same statistical modeling framework can be applied in this setting, with the outcome modeled as a binary preference. We use such a setup to illustrate how pairwise comparisons can be modeled and to examine length bias, which has often been observed in pairwise evaluation settings. Of course, length bias can also be captured in absolute score setups similarly to other biases in previous sections.
Let’s imagine that Florence wants to compare the quality of outputs generated by three different LLMs. She chooses a pairwise evaluation format, where each grader (e.g., herself or an autograder) is repeatedly shown two responses to the same prompt - each generated by a different LLM - and must select the better response. An example of such data can be seen in the left panel of fig. 7. Each bar represents a pairwise comparison (e.g., “LLM A vs. LLM B”), and its height reflects how frequently the first listed model (e.g., LLM A) was chosen. To model this, we can use a binomial GLM with a logit link function. The outcome variable $y_{i}$ indicates whether the first model in the pair was chosen ($y_{i}=1$) or not ($y_{i}=0$), and we include a categorical effect in the model to denote the LLM pair being compared.
@@
Florence’s younger brother, always up-to-date with ML controversies, recently told her that some autograders may systematically prefer longer outputs even if those outputs are not of higher quality, a phenomenon commonly referred to as length bias [Zheng et al., 2023; Dubois et al., 2024]. To capture this bias, she adds a continuous predictor capturing the token-length difference between the two outputs. As there are two graders (herself and the autograder), which might have different biases, she computes one such predictor per grader. To test the existence of the length bias formally, she compares the model with and without this term (cf. fig. 11 in appendix B). For demonstration purposes, let’s look at the model with grader-specific length bias here:
@@
$\begin{gathered}y_{i}\sim\text{Binomial}(1,p_{i})\\ \text{logit}(p_{i})=\beta_{0}+\beta_{1}(X_{i}^{\text{LLM pair}})+\beta_{2}(X_{i}^{\text{grader}})+\beta_{3}(X_{i}^{\text{grader}})\cdot X_{i}^{\text{lengthDiff}}\\ \beta_{3}(X_{i}^{\text{grader}})\sim\mathcal{N}(\mu_{\text{lengthDiff}},\sigma_{\text{lengthDiff}}^{2})\\ \mu_{\text{lengthDiff}}\sim\mathcal{N}(0,0.5)\\ \sigma_{\text{lengthDiff}}\sim\text{HalfNormal}(1.0)\\ \end{gathered}$ (7)
where $y_{i}$ is a binary outcome indicating whether the first-listed LLM was preferred. The intercept $\beta_{0}$ is the overall tendency to prefer the first-listed model, $\beta_{1}(X_{i}^{\text{LLM pair}})$ captures pair-specific preferences (e.g., “LLM A vs. LLM B”) and $\beta_{2}(X_{i}^{\text{grader}})$ captures each grader’s overall tendency to prefer the first-listed model. The predictor $X_{i}^{\text{lengthDiff}}$ is the token-length difference between the two responses. The grader-specific slope coefficient $\beta_{3}(X_{i}^{\text{grader}})$ quantifies how sensitive each grader is to length differences and is drawn from a hierarchical distribution with mean $\mu_{\text{lengthDiff}}$ and standard deviation $\sigma_{\text{lengthDiff}}$. Positive values of $\mu_{\text{lengthDiff}}$ indicate a preference for longer outputs. The hierarchical structure captures both the average length bias across graders and the variability among individual graders.
Figure 7: Illustration of how to use Bayesian GLMs to address Question 5 (Do autograders favour longer outputs?) using simulated data. Left panel: Proportion of pairwise preferences across three LLM pairs (A vs. B, A vs. C, B vs. C). Preference rate values represent the fraction of cases where the first listed model was selected over the second. Error bars represent 95% confidence intervals. Right panel: Posterior distributions of estimated effects from the GLM for pairwise comparisons (log-odds scale). Horizontal blue lines represent the 95% credible intervals, and the dashed red vertical line indicates a null effect ($\beta$=0). The pair effect terms represent the relative preference between specific pairs of LLMs, indicating which LLM is generally preferred. The grader length bias terms quantify each grader’s sensitivity to token-length differences when making choices. A positive length bias indicates a preference towards longer outputs.
@@
Importantly, once we have estimated the probability of choosing an LLM over another, we can compare these probabilities across pairs. This allows to identify rational (transitive) and irrational (intransitive) patterns of decision-making, such as cyclic dependencies (e.g., preferring A over B, B over C, but C over A). Such intransitivities exist in LLM evaluations [Xu et al., 2025]. Traditional models like the Bradley–Terry model implicitly assume transitivity and thus cannot capture these cycles. Recent approaches have proposed either removing intransitivities from datasets [Yu et al., 2025] or explicitly quantifying them [Zhang et al., 2025; Liu et al., 2024a; Zhao et al., 2024]. GLMs naturally capture these intransitivities alongside grader biases and differences in LLM performance.
@@
Inspecting the estimated effects in the right panel of fig. 7, we observe a positive effect for the grader-specific length bias parameter ($\beta_{3}$), particularly for the autograder. From this, Florence can infer that the autograder is more likely to select longer outputs, irrespective of their intrinsic quality. In other words, there appears to be an implicit association between output length and perceived correctness in the autograder’s judgments. By explicitly quantifying such biases within the model, Florence can more reliably interpret differences in LLM rankings. For example, if LLM A wins most comparisons but consistently produces longer outputs, Florence might question: “Is LLM A genuinely better, or simply more verbose? Can I really trust the autograder’s judgements?”. Additionally, by examining the estimated LLM pair parameters ($beta_{1}$), she can verify whether the observed preferences follow a consistent ranking or if there are intransitive (cyclical) patterns. Here, the estimated parameters indicate a consistent ordering: LLM A tends to be preferred to LLM B and LLM C, and LLM B tends to be preferred to LLM C. This integrated statistical framework empowers her to disentangle and quantify these systematic biases and assess preference consistency, leading to deeper and more reliable conclusions.
@@ Conclusion
In this paper, we introduced a statistical framework for evaluating autograders using Bayesian GLMs. By jointly modelling the evaluation outcome and the scoring process, this approach enables researchers to assess both LLM performance and autograder behaviour within a single analysis. Through a series of examples, we followed the journey of a fictional researcher toward evaluating autograders. We used simulated data throughout to explore a wide range of evaluation questions and settings, and to illustrate key modelling principles in a controlled and reproducible way.
Specifically, we showed how this framework can be used to compute richer, uncertainty-informed estimates of inter-rater agreement that, unlike traditional metrics, help us understand the sources of rater disagreement. We also demonstrated how it can quantify various types of biases (e.g., self-bias and length bias), capture individual-level differences both among graders and items, and improve the estimation of group-level trends through hierarchical modelling. Additionally, to highlight the method’s flexibility, we showed how it can easily be adapted to different evaluation formats, including absolute scoring and pairwise comparisons.
The examples presented are by no means exhaustive. Many other applications and extensions of GLMs are possible, and we hope this work provides a clear and practical starting point for researchers seeking to adapt the framework to their own evaluation scenarios. To support practical adoption, we have summarised common evaluation questions and their implementation in a GLM setup in table 1. All models presented in this paper are implemented in the open-source HiBayes package, and reproducible notebooks for all examples are available in a public repository.
@@ Appendix B Supplementary figures: Model comparisons
Figure 11: Model comparison for the statistical analysis in Question 5 (Do autograders favour longer outputs?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing these models tests whether including a length-effect term (eq. 7) significantly improves predictive performance. The model including the length effect clearly outperforms the simpler model without it, justifying a closer investigation into grader-specific length biases.
@@ Appendix C Supplementary question: Is the grading scale well calibrated?
Having an ordinal scale, instead of just a description, is of course very useful. But this scale is not as simple to interpret as an interval scale (e.g., temperature). If Florence wants to make claims like “autograder A increases the score by 1 point,” she needs to be aware of whether the intervals between categories are equivalent. If they are not (which is often the case with ordinal scales), a 1-point increase will mean something different depending on where it occurs on the scale. For example, moving from 5 to 6 might represent crossing some fundamental threshold, while moving from 9 to 10 might just be a small qualitative improvement between two already good responses.